20190108 DX ARGOS On-Sky Tests

05:40 Install old (feb 2018) FF matrix.
  • The DX AdSec configuration is back to its Pre shutdown/repair configuration
  • Closed ARGOS Loop
  • TT + LO gain is ok
  • HO gain of 0.05 causes stripes

06:00 put Act 573 and 450 in position do-not-use list
  • ARGOS Closed loop seem to have trouble with LO gain. See jumping more now than earlier in the evening. Saw this jumping in the past.

06:30 Put 450 back in use. It is on the bottom edge
  • Put 380 in position do not use list
  • Can now put ARGOS gains to nominal values. 1.2, 0.4, 0.3

07:00 Put 582 in position do-not-use list
  • more striping now
  • Is it the actuator or change in conditions

07:50 Put 582 back into use

-- DougMiller - 09 Jan 2019
Topic revision: r2 - 10 Jan 2019, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback